# Deep Learning for Cross-Domain Data Fusion in Urban Computing: Taxonomy, Advances, and Outlook

### v1

@@ 1 Introduction
Cities, indispensable components of modern civilization, have undergone transformative trajectories propelled by human advancements in cultural, financial, political, and technological domains Chen et al. (2023); Xu et al. (2023f); Zheng (2015). Despite their pivotal role in societal progress, the unprecedented surge in global urbanization since the 19th century has precipitated formidable sustainability challenges including energy consumption Wang et al. (2020a); Wu et al. (2019a), environmental pollution Liang et al. (2019a); Khan et al. (2022), socio-economic disparities Yao and Jiang (2021); Li et al. (2019), and urban traffic issues Kruszyna et al. (2021); Li et al. (2019). In the 21st century, the profound strides achieved in machine learning and spatio-tempral data mining have manifested in myriad successful applications across diverse domains, such as finance Huang et al. (2020); Ozbayoglu et al. (2020), biology Mahmud et al. (2021); Tang et al. (2019), and healthcare Stiglic et al. (2020); Shahid et al. (2019). This surge in technological prowess has incited a notable shift in research focus, with scholars now directing their attention toward harnessing these advancements to optimize the intricate facets of urban planning, operations, management, etc. A pivotal contribution to this evolving discourse is the pioneering work Zheng et al. (2014a), which encapsulated and elucidated these endeavors by introducing the concept of Urban Computing. This paradigm leverages sensing technologies and expansive computing infrastructure to scrutinize voluminous data emanating from urban spaces. The fundamental objective is to gain profound insights into the dynamics of cities, thereby addressing challenges such as traffic congestion Ji et al. (2023a); Zhang et al. (2021c); Pan et al. (2020), energy consumption Johari et al. (2020); Hashem et al. (2023), and air quality pollution Liang et al. (2023); Yi et al. (2018, 2020).
@@
Urban computing necessitates the integration of extensive and diverse datasets sourced from various sources and modalities Zheng et al. (2014a); Zheng (2015), also referred to as Cross-Domain Data Fusion, which arises from the recognition that relying solely on a singular data source or modality may prove inadequate for the holistic implementation of urban tasks. In recent years, a growing number of studies in urban computing are expanding cross-domain data fusion to encompass diverse sources like sensors Ji et al. (2016), satellites Liu et al. (2023g), social media Zhao et al. (2020), and citizen-generated data Zhao et al. (2016b). Additionally, there is a trend towards introducing new data modalities, including text (e.g., social media posts Zhao et al. (2016b) and geographic information Huang et al. (2022a); Chen et al. (2021c)) and images (e.g., satellite Liu et al. (2023g); Xi et al. (2022); Li et al. (2022b) and street-view images Liu et al. (2023g); Li et al. (2022b); Huang et al. (2023a)). Figure 1 depicts the cross-domain data fusion in urban computing from the views of data modality and source.
@@
Prior research, exemplified by Zheng et al. (2014a), emphasized the critical role of cross-domain data fusion in amalgamating information from multiple sources. With the emergence of data fusion studies in urban computing, Zheng (2015) classified the related fusion methodologies into three types: stage-based, feature level-based, and semantic-based data fusion. Furthermore, within the purview of semantic-based methods, a more intricate taxonomy emerges, delineating four subtypes: multi-view learning-based, similarity-based, probabilistic dependency-based, and transfer learning-based methods.
@@
Over the past decade, the emergence of Deep Learning (DL) has asserted dominance in processing spatio-temporal data in urban computing Wang et al. (2020b); Jin et al. (2023a). In contrast to traditional machine learning methods, these models present distinct superiority, characterized by larger model capacity, automated feature extraction capabilities, and inherent compatibility with cross-domain data fusion. The paradigm shift derived from deep learning renders previous surveys, especially Zheng (2015), on urban data fusion somewhat obsolete, as traditional taxonomy may not aptly capture the nuances and differences among these advanced methodologies. In light of this issue, our survey is dedicated to bridging this gap and provides a contemporary perspective by offering a comprehensive and updated taxonomy that aligns with the era of deep learning. Through a thorough examination of deep learning-based cross-domain data fusion methods, we seek to establish a robust foundation for understanding and navigating the landscape of urban computing.
@@
Specifically, we commence by presenting a novel taxonomy that classifies existing urban data sources into five distinct types, while concurrently categorizing fusion methods into four types. This systematic classification offers valuable insights into the intricate integration of diverse modalities within urban computing. Secondly, we systematically categorize widely used datasets and outline common application scenarios for urban computing fusion models. Thirdly, inspired by the breakthrough of Large Language Models (LLMs) in financial time series Wu et al. (2023c); Yu et al. (2023), medicine Thirunavukarasu et al. (2023); Singhal et al. (2023), law Cui et al. (2023), and climate Schimanski et al. (2023); Laud et al. (2023), we further summarize inspiring research that integrates LLMs into urban computing, which complements the basic taxonomy of cross-domain data fusion in urban computing.
@@
Related Surveys. Several surveys have recently explored the application of deep learning-based data fusion across various domains. Notably, Zheng (2015) conducted a comprehensive investigation into the methodologies proposed for cross-domain big data fusion prior to 2015. Their research reveals that machine learning-based data fusion was the de-facto dominant approach in urban computing during that period, albeit with challenges in comprehending cross-modal relations. Subsequently, with the remarkable success of deep learning models such as recurrent neural networks (RNN) and convolutional neural networks (CNN) in representation learning, Wang et al. (2020b) conducted an exhaustive review on the utilization of deep learning for spatio-temporal data mining, with a special focus on the fusion of multi-source spatio-temporal data. Liu et al. (2020a) further provided a summary of deep learning-based data fusion methodologies for urban big data fusion before 2020. Though these surveys shed light on the advancements in deep learning to their respective fields, they primarily concentrated on specific aspects and do not extensively cover cross-modal data fusion or the latest paradigms, including contrast-based and generation-based approaches which have gained significant popularity since 2021. In other words, none of them have provided a comprehensive and up-to-date taxonomy framework for data fusion methodologies in the context of urban computing.
@@
Furthermore, existing surveys lack a specific focus on the data and application perspectives of these DL models within the field of urban computing. For instance, Gao et al. (2022b) summarized the fusion models for spatio-temporal data based on generative adversarial networks (GAN), while Deldari et al. (2022) concentrated on self-supervised representation learning on the fusion of multi-modal data in the general domain. Besides, there are a couple of surveys investigating the deep learning-based data fusion in specific applications (i.e., a subdomain) of urban computing, such as crowd flow prediction Xie et al. (2020), intelligent transportation Yuan and Li (2021), and social event detection Afyouni et al. (2022). These facts highlight the need for a new survey to serve as a guide for future endeavors in data fusion in urban computing concepts.
@@
Zheng (2015) 2015 IEEE Trans. Big Data
Wang et al. (2020b) 2020 IEEE TKDE
Liu et al. (2020a) 2020 Information Fusion
Xie et al. (2020) 2020 Information Fusion
Yuan and Li (2021) 2021 Springer Data Sci. Eng.
Afyouni et al. (2022) 2022 Information Fusion
Gao et al. (2022b) 2022 ACM TIST
Deldari et al. (2022) 2022 Unpublished
@@
  • Future Research Outlook. We endeavor to identify and provide detailed explanations of several promising directions for future research, covering various aspects, including data privacy protection, the establishment of open benchmarks, the diversification of applications, and the optimization of efficiency. Moreover, capitalizing on the progress of LLMs (e.g., GPT-4 OpenAI (2023) and Sora Brooks et al. (2024)) and their notable benefits in fusing multi-modal and multi-source data, we further investigate their innovative applications and propose potential approaches for urban computing.
@@ 2 Taxonomy
In Section 3, from the data perspective, we divided the data normally utilized in urban computing into five categories according to data sources: geographical data, traffic data, social media data, demographic data and environmental data. Additionally, we have also classified the data from a modality perspective, including spatio-temporal data, visual data, textual data, and other types of data. These two categorizations allows for a systematic understanding and analysis of the different types of data used in urban computing research. We further present a comprehensive overview of public datasets used in cross-domain data fusion in urban computing in Table 2.
@@ 3.1 Overview
Based on the characteristics of the data and their diverse sources from various domains, this survey categorizes datasets utilized in the field of cross-domain data fusion in urban computing into six segments, including geographical data, traffic data, social network data, demographic data, environment data, and other data (i.e., data that cannot be categorized into the aforementioned types, such as healthcare data). As illustrated in Figure 1 and 2, these categories are defined as follows:
@@
ArcGIS https://developers.arcgis.com Liu et al. (2023g)
PlanetScope https://developers.planet.com/docs/data/planetscope/ Li et al. (2022b)
@@
Jenkins et al. (2019)
OpenStreetMap https://www.openstreetmap.org/ Yin et al. (2023b)
@@
Liu et al. (2023g); Alfarrarjeh et al. (2021)
Street-View Image Image Tencent Map https://lbs.qq.com/tool/streetview/index.html Huang et al. (2023a)
@@
Xi et al. (2022); Ruan et al. (2022)
WeChat POIs https://open.weixin.qq.com Wang et al. (2021c)
@@
Bing et al. (2023); Zhao et al. (2016a); Balsebre et al. (2022); Chen et al. (2015); Huang and Wang (2016); Jenkins et al. (2019)
Wikipedia POIs https://www.wikipedia.org Zhao et al. (2017)
AMap Service https://lbs.amap.com Bai et al. (2023)
@@
Chandra et al. (2021); Keerthi Chandra et al. (2020); Gao et al. (2023b)
Flickr POIs https://www.flickr.com/services/developer/api/ He et al. (2019)
@@
Chen et al. (2022b)
Shenzhou UCar https://bit.ly/2MG47xz Guo et al. (2022)
@@
Wang et al. (2021a, 2023c); Jenkins et al. (2019)
VED https://github.com/gsoh/VED Oh et al. (2019); Zhang et al. (2022c)
@@
Yuan et al. (2011, 2010); Pan et al. (2019); Lv et al. (2018)
DiDi Traffic https://outreach.didichuxing.com/research/opendata/ Yuan et al. (2021a); Luo et al. (2021a); Rajeh et al. (2023); Yao et al. (2022); Tedjopurnomo et al. (2021)
@@
Yin et al. (2023b, 2020)
California-PEMS http://pems.dot.ca.gov Arslan Ay et al. (2009); Sun et al. (2023)
METR-LA https://www.metro.net Lablack and Shen (2023); Liu et al. (2023a)
Large-ST https://github.com/liuxu77/LargeST Liu et al. (2023f)
MobileBJ https://github.com/FIBLAB/DeepSTN/issues/4 Lin et al. (2019); Keerthi Chandra et al. (2020); Chandra et al. (2021)
TaxiBJ https://goo.su/aQyjTAz Liang et al. (2021); Bai et al. (2019a); Qu et al. (2022); Ji et al. (2023b); Zhang et al. (2021b); Fu et al. (2019)
Taffic Flow Spatio-temporal Graph BikeNYC https://citibikenyc.com/ Lin et al. (2019); Bai et al. (2019a); Qu et al. (2022); Ji et al. (2023b)
Road Network Spatial Graph OpenStreetMap https://www.openstreetmap.org Yin et al. (2020); Balsebre et al. (2022); Luo et al. (2021a); Yuan et al. (2021a); Geng et al. (2019b)
US Census Bureau https://www.census.gov/data.html Zhang et al. (2023c)
Traffic Data LaDe https://cainiaotechai.github.io/LaDe-website/ Wu et al. (2023b)
Logistics Spatio-temporal Trajectory JD Logistics https://corporate.jd.com/ourBusiness#jdLogistics Ruan et al. (2022)
@@
Bing et al. (2023); Zhao et al. (2016a, b); Yuan et al. (2017); Vu et al. (2016); Wu et al. (2015); Shen et al. (2018)
Wang et al. (2017b, 2016a, 2016b, 2015); Miyazawa et al. (2019); Liu et al. (2022c)
Common Crawl https://registry.opendata.aws/commoncrawl/ Zhao et al. (2020)
@@
Zhao et al. (2020, 2016b)
Text Text Weibo Traffic Police http://open.weibo.com/developers/ You et al. (2022)
YFCC100M https://goo.su/jzaDU Zhao et al. (2017); Yin et al. (2021); He et al. (2019)
NUS-WIDE https://goo.su/dWPQZcD Yin et al. (2021, 2019)
Geo-tagged Image & Video Image&Video GeoUGV https://qualinet.github.io/databases/video/ Lu et al. (2016)
Jiepang User Check-in https://jiepang.app/ Fu et al. (2019)
@@
Chen et al. (2015); Yuan et al. (2017)
Social Media Data Users’ Info Time Series WeChat Mobility https://open.weixin.qq.com/ Wang et al. (2021c)
Crime Time Series NYC Crime https://opendata.cityofnewyork.us/ Zhang et al. (2021b)
Land Use SG https://www.ura.gov.sg/Corporate/Planning/Master-Plan Li et al. (2023d)
Land Use Time Series Land Use NYC https://goo.su/puTuG Li et al. (2023d)
@@
Xi et al. (2022); Li et al. (2022b); Bai et al. (2023)
TipDM China Weather https://www.tipdm.org/ Liu et al. (2022b)
DarkSky Weather https://support.apple.com/en-us/102594 Yuan et al. (2021a)
WeatherNY https://opendata.cityofnewyork.us/ Wang et al. (2021a)
WeatherChicago https://data.cityofchicago.org/ Wang et al. (2021a)
Weather Underground https://www.wunderground.com/ You et al. (2022)
DidiSY https://www.didiglobal.com/ Bai et al. (2019b)
WD_BJ weather https://goo.su/DmHFHd Ma et al. (2023a)
Meteorology Time Series WD_USA weather https://goo.su/RVhBA Ma et al. (2023a)
Greenery Time Series Google Earth https://earth.google.com/ You et al. (2022)
Environment Data Air Quality Time Series UrbanAir https://goo.su/hfzNB53 Zheng et al. (2015, 2013, 2014b)
KnowAir https://github.com/shuowang-ai/PM2.5-GNN Wang et al. (2020c); Yu et al. (2021); Zhang et al. (2022b); Xu et al. (2023c)

### v2

@@ 1 Introduction
Cities, indispensable components of modern civilization, have undergone transformative trajectories propelled by human advancements in cultural, financial, political, and technological domains [41, 321, 390, 300]. Despite their pivotal role in societal progress, the unprecedented surge in global urbanization since the 19th century has precipitated formidable sustainability challenges including energy consumption [281, 303], environmental pollution [161, 135, 274], socio-economic disparities [330, 150], and urban traffic issues [140, 150, 212, 167]. In the 21st century, the profound strides achieved in machine learning and spatio-tempral data mining have manifested in myriad successful applications across diverse domains, such as finance [109, 215, 19], biology [194, 257], and healthcare [251, 238]. This surge in technological prowess has incited a notable shift in research focus, with scholars now directing their attention toward harnessing these advancements to optimize the intricate facets of urban planning, operations, management, etc. A pivotal contribution to this evolving discourse is the pioneering work [391], which encapsulated and elucidated these endeavors by introducing the concept of Urban Computing. This paradigm leverages sensing technologies and expansive computing infrastructure to scrutinize voluminous data emanating from urban spaces. The fundamental objective is to gain profound insights into the dynamics of cities, thereby addressing challenges such as traffic congestion [119, 375, 218], energy consumption [128, 98, 223], and air quality pollution [166, 335, 334].
@@
Urban computing necessitates the integration of extensive and diverse datasets sourced from various sources and modalities [391, 390, 67], also referred to as Cross-Domain Data Fusion, which arises from the recognition that relying solely on a singular data source or modality may prove inadequate for the holistic implementation of urban tasks. For example, in the realm of traffic prediction, it becomes imperative to assimilate meteorological forecast data with geographical information. This involves taking into account the congestion induced by rainfall and the influence of school and business hours on traffic flow during peak periods. In the context of urban planning, one must combine population density and economic activity data. This includes evaluating factors such as population density and income levels when devising plans for new commercial districts to ascertain their viability. Furthermore, in the field of public safety management, integrating crime data with socio-economic data is crucial. This entails considering unemployment rates and education levels in high-crime areas when strategizing police deployment plans. In recent years, a growing number of studies in urban computing are expanding cross-domain data fusion to encompass diverse sources like sensors [121], satellites [186], social media [380], and citizen-generated data [383]. Additionally, there is a trend towards introducing new data modalities, including text (e.g., social media posts [383] and geographic information [110, 43]) and images (e.g., satellite [186, 309, 154] and street-view images [186, 154, 112]). Figure 1 depicts the cross-domain data fusion in urban computing from the views of data modality and source.
@@
Prior research, exemplified by Zheng et al. [391], emphasized the critical role of cross-domain data fusion in amalgamating information from multiple sources. With the emergence of data fusion studies in urban computing, Zheng [390] classified the related fusion methodologies into three types: stage-based, feature level-based, and semantic-based data fusion. Furthermore, within the purview of semantic-based methods, a more intricate taxonomy emerges, delineating four subtypes: multi-view learning-based, similarity-based, probabilistic dependency-based, and transfer learning-based methods.
Figure 2: Times cited and publications over time for Deep Learning in Urban Computing on prestigious venues(Source: Web of Science)
@@
Over the past decade, the emergence of Deep Learning (DL) has asserted dominance in processing spatio-temporal data in urban computing [282, 125]. In contrast to traditional machine learning methods, these models present distinct superiority, characterized by larger model capacity, automated feature extraction capabilities, and inherent compatibility with cross-domain data fusion. As Figure 2 illustrates, there has been a significant uptick in both the number of papers published and the citations received for research related to deep learning in urban computing since 2015. The paradigm shift derived from deep learning renders previous surveys, especially [390], on urban data fusion somewhat obsolete, as traditional taxonomy may not aptly capture the nuances and differences among these advanced methodologies. In light of this issue, our survey is dedicated to bridging this gap and provides a contemporary perspective by offering a comprehensive and updated taxonomy that aligns with the era of deep learning. Through a thorough examination of deep learning-based cross-domain data fusion methods, we seek to establish a robust foundation for understanding and navigating the landscape of urban computing.
@@
Specifically, we commence by presenting a novel taxonomy that classifies existing urban data sources into five distinct types, while concurrently categorizing fusion methods into four types. This systematic classification offers valuable insights into the intricate integration of diverse modalities within urban computing. Secondly, we systematically categorize widely used datasets and outline common application scenarios for urban computing fusion models. Thirdly, inspired by the breakthrough of Large Language Models (LLMs) in financial time series [306, 345], medicine [262, 244], law [51], and climate [236, 146], we further summarize inspiring research that integrates LLMs into urban computing, which complements the basic taxonomy of cross-domain data fusion in urban computing.
@@
Related Surveys. Several surveys have recently explored the application of deep learning-based data fusion across various domains. Notably, Zheng [390] conducted a comprehensive investigation into the methodologies proposed for cross-domain big data fusion prior to 2015. Their research reveals that machine learning-based data fusion was the de-facto dominant approach in urban computing during that period, albeit with challenges in comprehending cross-modal relations. Subsequently, with the remarkable success of deep learning models such as recurrent neural networks (RNN) and convolutional neural networks (CNN) in representation learning, Wang et al. [282] conducted an exhaustive review on the utilization of deep learning for spatio-temporal data mining, with a special focus on the fusion of multi-source spatio-temporal data. Liu et al. [177] further provided a summary of deep learning-based data fusion methodologies for urban big data fusion before 2020. Though these surveys shed light on the advancements in deep learning to their respective fields, they primarily concentrated on specific aspects and do not extensively cover cross-modal data fusion or the latest paradigms, including contrast-based and generation-based approaches which have gained significant popularity since 2021. In other words, none of them have provided a comprehensive and up-to-date taxonomy framework for data fusion methodologies in the context of urban computing.
@@
Furthermore, existing surveys lack a specific focus on the data and application perspectives of these DL models within the field of urban computing. For instance, Gao et al. [76] summarized the fusion models for spatio-temporal data based on generative adversarial networks (GAN), while Deldari et al. [56] concentrated on self-supervised representation learning on the fusion of multi-modal data in the general domain. Besides, there are a couple of surveys investigating the deep learning-based data fusion in specific applications (i.e., a subdomain) of urban computing, such as crowd flow prediction [315], intelligent transportation [348], and social event detection [2]. These facts highlight the need for a new survey to serve as a guide for future endeavors in data fusion in urban computing concepts.
@@
Zheng [390] 2015 IEEE Trans. Big Data
Wang et al. [282] 2020 IEEE TKDE
Liu et al. [177] 2020 Information Fusion
Xie et al. [315] 2020 Information Fusion
Yuan and Li [348] 2021 Springer Data Sci. Eng.
Afyouni et al. [2] 2022 Information Fusion
Gao et al. [76] 2022 ACM TIST
Deldari et al. [56] 2022 Unpublished
@@
  • Future Research Outlook. We endeavor to identify and provide detailed explanations of several promising directions for future research, covering various aspects, including data privacy protection, the establishment of open benchmarks, the diversification of applications, and the optimization of efficiency. Moreover, capitalizing on the progress of LLMs (e.g., GPT-4 [211] and Sora [22]) and their notable benefits in fusing multi-modal and multi-source data, we further investigate their innovative applications and propose potential approaches for urban computing.
@@ 2 Taxonomy
In Section 3, from the data perspective, we divided the data normally utilized in urban computing into five categories according to data sources: geographical data, traffic data, social media data, demographic data and environmental data. Additionally, we have also classified the data from a modality perspective, including spatio-temporal data, visual data, textual data, and other types of data. These two categorizations allow for a systematic understanding and analysis of the different types of data used in urban computing research. We further present a comprehensive overview of public datasets used in cross-domain data fusion in urban computing in Table 2.
@@ 3.1 Overview
Based on the characteristics of the data and their diverse sources from various domains, this survey categorizes datasets utilized in the field of cross-domain data fusion in urban computing into six segments, including geographical data, traffic data, social network data, demographic data, environment data, and other data (i.e., data that cannot be categorized into the aforementioned types, such as healthcare data). As illustrated in Figure 1 and 3, these categories are defined as follows:
@@
ArcGIS https://developers.arcgis.com [186]
PlanetScope https://developers.planet.com/docs/data/planetscope/ [154]
@@
[116]
OpenStreetMap https://www.openstreetmap.org/ [337]
@@
[186, 4]
Street-View Image Image Tencent Map https://lbs.qq.com/tool/streetview/index.html [112]
@@
[309, 235]
WeChat POIs https://open.weixin.qq.com [277]
@@
[20, 381, 13, 42, 107, 116]
Wikipedia POIs https://www.wikipedia.org [386]
AMap Service https://lbs.amap.com [10]
@@
[33, 134, 77]
Flickr POIs https://www.flickr.com/services/developer/api/ [99]
@@
[37]
Shenzhou UCar https://bit.ly/2MG47xz [93]
@@
[272, 288, 116]
VED https://github.com/gsoh/VED [209, 372]
@@
[350, 351, 217, 191]
DiDi Traffic https://outreach.didichuxing.com/research/opendata/ [349, 188, 228, 328, 261]
@@
[337, 339]
California-PEMS http://pems.dot.ca.gov [9, 254]
METR-LA https://www.metro.net [143, 171]
Large-ST https://github.com/liuxu77/LargeST [182]
MobileBJ https://github.com/FIBLAB/DeepSTN/issues/4 [170, 134, 33]
TaxiBJ https://goo.su/aQyjTAz [164, 11, 226, 120, 368, 74]
Taffic Flow Spatio-temporal Graph BikeNYC https://citibikenyc.com/ [170, 11, 226, 120]
Road Network Spatial Graph OpenStreetMap https://www.openstreetmap.org [339, 13, 188, 349, 84]
US Census Bureau https://www.census.gov/data.html [366]
Traffic Data LaDe https://cainiaotechai.github.io/LaDe-website/ [305]
Logistics Spatio-temporal Trajectory JD Logistics https://corporate.jd.com/ourBusiness#jdLogistics [235]
@@
[20, 381, 383, 352, 270, 301, 240]
[289, 283, 285, 284, 200, 184]
Common Crawl https://registry.opendata.aws/commoncrawl/ [380]
@@
[380, 383]
Text Text Weibo Traffic Police http://open.weibo.com/developers/ [342]
YFCC100M https://goo.su/jzaDU [386, 340, 99]
NUS-WIDE https://goo.su/dWPQZcD [340, 338]
Geo-tagged Image & Video Image&Video GeoUGV https://qualinet.github.io/databases/video/ [187]
Jiepang User Check-in https://jiepang.app/ [74]
@@
[42, 352]
Social Media Data Users’ Info Time Series WeChat Mobility https://open.weixin.qq.com/ [277]
Crime Time Series NYC Crime https://opendata.cityofnewyork.us/ [368]
Land Use SG https://www.ura.gov.sg/Corporate/Planning/Master-Plan [156]
Land Use Time Series Land Use NYC https://goo.su/puTuG [156]
@@
[309, 154, 10]
TipDM China Weather https://www.tipdm.org/ [178]
DarkSky Weather https://support.apple.com/en-us/102594 [349]
WeatherNY https://opendata.cityofnewyork.us/ [272]
WeatherChicago https://data.cityofchicago.org/ [272]
Weather Underground https://www.wunderground.com/ [342]
DidiSY https://www.didiglobal.com/ [12]
WD_BJ weather https://goo.su/DmHFHd [192]
Meteorology Time Series WD_USA weather https://goo.su/RVhBA [192]
Greenery Time Series Google Earth https://earth.google.com/ [342]
Environment Data Air Quality Time Series UrbanAir https://goo.su/hfzNB53 [399, 396, 392]
KnowAir https://github.com/shuowang-ai/PM2.5-GNN [286, 346, 370, 318]